# Attention-Guided Perturbation for Unsupervised Image Anomaly Detection

### v1

@@ III-D Anomaly Map
IV Experiment
IV-A Experimental Setups
Datasets. MVTec-AD [1] is one of the most widely used industrial anomaly detection datasets. It has a training set consisting of 3629 normal images and a test set of 467/1,258 normal/anomaly images. All the images are divided into 15 categories. In this paper, we conduct experiments with two settings. One involves training a single model for all classes (multi-class setting), while the other involves training a separate model for each class (one-class setting). The former is a more efficient yet challenging task.
VisA [42] is another challenging anomaly detection dataset. It has a total of 10,821 images divided into 12 different categories. The training set has 8,659 normal images, and the test set includes 962/1,200 normal/anomaly images.
MVTec-3D [43] is a collection of 4,147 scans obtained through the industrial 3D sensor. It covers 10 different categories, each accompanied by both RGB images and corresponding 3D point cloud data. The training set consists of 2,656 images that include only anomaly-free samples. The test set comprises 1,197 images, which encompass both normal and anomalous samples. In our experiments conducted, only the RGB images were utilized, with the 3D point cloud data not being considered.
Evaluation Metrics. Referring to prior research [5, 3, 2], we report the anomaly detection and localization performance with Image-level and Pixel-level Area Under the Receiver Operating Curve, denoted by I-AUC and P-AUC respectively. We also adopt Per-Region-Overlap[44] to better evaluate model’s capability, denoted by PRO.
Implement Details. Each image is resized to $224\times 224$. The feature extractor is initialized by pre trained DINO(ViT-S-16) [39], which is frozen during training. We employ AdamW[45] optimizer with a weight decay of $1\times 10^{-4}$. We train our framework for 500 epochs on a single GPU (NVIDIA RTX 4090) with a batch size of 32. The learning rate is initially set to $1\times 10^{-3}$ and reduced by a factor of 0.1 after 200 epochs. The ratio of added noise at the image level linearly increases from 0.6 to 1.0 from the 100th to 400th epochs. The intensity of added noise $\alpha$ at the feature level linearly increases from 0 to 1.0 from the 1st to 400th epochs. The parameter $\eta$ in EMA is set to 0.9999, and the mean teacher is updated every 10 steps. In particular, we performed a 32-fold data augmentation (rotate and flip) under a few-shot setting.
TABLE I: Anomaly detection results on MVTec AD under multi-class setting. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
@@ IV-A Experimental Setups
Bottle 97.5 / 87.6 / 80.7 97.7 / 91.2 / 90.6 99.7 / 98.1 / 93.2 99.7 / 98.4 / - 100.0 / 98.8 / 94.3 100.0 / 99.1 / 95.0
Cable 57.8 / 71.3 / 40.1 87.6 / 88.1 / 85.4 95.2 / 97.3  85.7 94.8 / 96.8 / - 98.5 / 96.2 / 90.2 99.7 / 98.0 / 92.9
Capsule 65.3 / 50.5 / 27.3 78.3 / 89.7 / 84.5 86.9 / 98.5  89.7 89.0 / 97.1 / - 95.4 / 98.3 / 92.0 93.8 / 98.5 / 91.9
Hazelnut 93.7 / 96.9 / 78.7 99.2 / 95.7 / 87.4 99.8 / 98.1 / 92.9 99.5 / 98.3 / - 99.8 / 99.0 / 95.2 100.0 / 98.9 / 93.6
Metal_nut 72.8 / 62.2 / 66.4 85.1 / 90.9 / 85.2 99.2 / 94.8 / 84.7 99.1 / 97.3 / - 99.7 / 96.4 / 92.4 100.0 / 95.9 / 92.6
Pill 82.2 / 94.4 / 53.9 78.3 / 89.7 / 81.9 93.7 / 95.0 / 94.7 95.7 / 95.7 / - 96.2 / 98.7 / 95.3 97.6 / 98.2 / 96.2
Screw 92.0 / 95.5 / 55.2 45.5 / 93.7 / 84.0 87.5 / 98.3 / 94.9 90.7 / 97.9 / - 91.3 / 99.0 / 93.5 91.8 / 99.0 / 95.5
Toothbrush 90.6 / 97.7 / 68.9 94.7 / 97.5 / 87.4 94.2 / 98.4 / 88.0 99.7 / 99.0 / - 98.9 / 99.1  90.9 100.0 / 98.9 / 91.2
Transistor 74.8 / 64.5 / 39.0 82.0 / 86.0 / 83.2 99.8 / 97.9 / 93.3 99.8 / 95.1 / - 98.8 / 93.9 / 76.8 98.7 / 93.8 / 84.3
Zipper 98.8 / 98.3 / 91.9 99.1 / 97.0 / 90.7 95.8 / 96.8 / 91.6 95.1 / 96.2 / - 97.6 / 95.9 / 87.2 99.7 / 97.6 / 90.5
Carpet 98.0 / 98.6 / 93.1 95.9 / 92.4 / 90.6 99.8 / 98.5 / 93.6 99.4 / 98.6 / - 99.5 / 99.0 / 94.7 100.0 / 99.2 / 95.6
Grid 99.3 / 98.7 / 92.1 49.8 / 46.7 / 88.6 98.2 / 96.5 / 91.1 98.5 / 96.6 / - 99.7 / 98.6 / 95.8 99.9 / 98.9 / 96.3
Leather 98.7 / 97.3 / 88.5 93.9 / 96.9 / 92.7 100.0 / 98.8 / 96.5 99.8 / 98.8 / - 100.0 / 99.6 / 97.9 100.0 / 97.5 / 97.6
Tile 99.8 / 98.0 / 97.0 93.7 / 93.1 / 90.6 99.3 / 91.8 / 78.6 96.8 / 92.4 / - 100.0 / 96.6 / 87.0 100.0 / 97.4 / 90.2
Wood 99.8 / 96.0 / 94.2 95.2 / 84.8 / 76.3 98.6 / 93.2 / 85.0 99.7 / 93.3 / - 98.7 / 96.4 / 88.0 98.9 / 96.6 / 90.2
Mean 88.1 / 87.2 / 71.1 85.1 / 88.9 / 86.5 96.5 / 96.8 / 90.2 97.2 / 96.8 / 90.7 98.3 / 97.7 / 91.4 98.7 / 98.0 / 92.9
@@
IV-B Anomaly Detection on MVTec-AD
@@
Baseline. We comprehensively evaluated representative methods in both one-class and multi-class settings. In the multi-class setting, we first compare our method with two representative one-class approaches: DRAEM[3] and SimpleNet[46], and then compare with methods specifically designed for the multi-class setting, including UniAD[5], RD4AD[8], DiAD[17], ViTAD[47]. Note that all methods are trained under the multi-class scenario. In the one-class setting, the compared models includes: RD4AD[8], PatchCore[28], DRAEM[3], DeSTSeg[7], Simplenet[46]. In the few-shot setting, We compared our method with some competitive methods and the models designed specifically for few-shot setting, including SPADE[48], PaDiM [2], PatchCore [28], WinCLIP+ [49], RWDA [50], FastRcon [51], PromptAD [52].
@@
Evaluation on MVTec-AD under the multi-class setting. As shown in  Table I, we conducted comprehensive comparative experiments on the MVTec-AD to demonstrate the superiority of our method. It can be observed that applying one-class methods to the multi-class scenario reflects undesirable performance. Furthermore, compared with existing multi-class methods, we outperform all of them, achieving the best 98.7% and 98.0% I-AUC and P-AUC, respectively. ViTAD is a competitive approach, while we outperform it in 12 out of 15 categories in detection capacity, and exceed it in 10 categories for localization performance. In the more challenging PRO metric, our model also achieves the best performance 92.9%, with a 1.5% improvement over ViTAD.  Fig. 3 shows qualitative comparisons between DiAD, UniAD, and our method, indicating that we achieve the most precise anomaly localization results across diverse anomalies.
Evaluation on MVTec-AD under the one-class setting. As shown in  Table II, we can observe that compared with the two synthetic data-based methods, DRAEM [3] and DeSTSeg [28], we outperform them by a large margin. Specifically, compared with DRAEM, we outperform it by 1.2% and 1.0% in I-AUC and P-AUC, respectively. And compared to DeSTSeg, we also bring gains of 0.6% I-AUC and 0.4% P-AUC. Moreover, we achieved the best localization result of 98.3% P-AUC, while the detection performance is 0.4% lower than SimpleNet, ranking second among previous arts. We argue that although SimpleNet performs well in the one-class setting, its performance significantly drops when adapted to the multi-class scenario (as shown in  Table I). Therefore, overall, our model maintains satisfactory performance across both settings.
Evaluation on MVTec-AD under the few-shot setting. As shown in  Table IV, to further demonstrate the generality and potential of our framework, we conduct experiments on MVTec-AD under the challenging few-shot anomaly detection setups. The experimental results demonstrate the superiority of our method, especially at pixel level. Specifically, our method achieves the best localization performance of 96.2% and 97.3% P-AUC under 2-shot and 4-shot settings. Compared to the competitive method PromptAD, we achieved a 0.8% P-AUC increase under 4-shot settings.Patchcore is a competitive anomaly detection model under the traditional task settings. However, when faced with the more challenging Few-shot setup, our approach outperforms Patchcore with improvements of 6.4%, 7.0%, and 9.0% in I-AUC, and 3.0%, 2.9%, and 3.6% in P-AUC under the 4-shot, 2-shot, and 1-shot setting, respectively.In summary, the results corroborate the method’s generality, particularly its competitive performance demonstrated across various task setups.
@@
TABLE II: Anomaly detection results on MVTec AD under one-class setting. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
@@
IV-C Anomaly Detection on VisA
Baseline. We further conduct experiments on the challenging VisA dataset and compare with state-of-the-art models including DRAEM [3], UniAD [5], Simplenet [46], DiAD [17], and ViTAD [47].
Evaluation on VisA. As shown in  Table III, our method achieves state-of-the-art (SOTA) performance. Compared with UniAD, we outperform it by 6.8% and 3.5% in I-AUC and P-AUC. Specifically, Faced with a competitive model like ViTAD, we outperform it in 10 out of 12 categories for detection and exceed it in 6 categories for localization. Compared with the previous state-of-the-art model (ViTAD), we improve upon it by 1.9%, 0.2% and 0.4% in I-AUC, P-AUC and PRO, respectively. It once again proves the superiority of our method.
IV-D Anomaly Detection on MVTec-3D
Baseline. To further verify the superiority of the proposed method, we conduct experiments on the MVTec-3D dataset and compare it with state-of-the-art models including DRAEM [3], UniAD [5], and DiAD [17].
Evaluation on MVTec-3D. As illustrated in Table V, our approach outperforms all others across every metric evaluated. Notably, when compared to UniAD, our model demonstrates a considerable advantage, with improvements of 6.0%, 1.5%, and 4.8% in the I-AUC, P-AUC, and PRO, respectively. Compared to the recent diffusion-based method DiAD, our approach has achieved significant improvements in the evaluation metrics, with enhancements of 0.3%, 1.6%, and 5.1% in the I-AUC, P-AUC, and PRO, respectively. These results substantiate the superior performance of our method.
TABLE III: Anomaly detection results on VisA. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
@@ IV-D Anomaly Detection on MVTec-3D
TABLE IV: Anomaly detection results on MVTec under few shot setting. Bold and underline indicate the best and the second best, respectively.

### v2

@@ 3.3 Anomaly Map
4 Experiment
4.1 Experimental Setups
Datasets. MVTec-AD Bergmann et al. (2019) is one of the most widely used industrial anomaly detection datasets. It has a training set consisting of 3629 normal images and a test set of 467/1,258 normal/anomaly images. All the images are divided into 15 categories. In this paper, we conduct experiments with two settings. One involves training a single model for all classes (multi-class setting), while the other involves training a separate model for each class (one-class setting). The former is a more efficient yet challenging task.
VisA Zou et al. (2022) is another challenging anomaly detection dataset. It has a total of 10,821 images divided into 12 different categories. The training set has 8,659 normal images, and the test set includes 962/1,200 normal/anomaly images, providing a comprehensive benchmark for evaluating detection performance.
MVTec-3D Bergmann et al. (2021) is a collection of 4,147 scans obtained through the industrial 3D sensor. It covers 10 different categories, each accompanied by both RGB images and corresponding 3D point cloud data. The training set consists of 2,656 images that include only anomaly-free samples. The test set comprises 1,197 images, which encompass both normal and anomalous samples. In our experiments conducted, only the RGB images were utilized, with the 3D point cloud data not being considered.
Evaluation Metrics. Referring to prior research You et al. (2022); Zavrtanik et al. (2021); Defard et al. (2021), we report the anomaly detection and localization performance with Image-level and Pixel-level Area Under the Receiver Operating Curve, denoted by I-AUC and P-AUC respectively. We also adopt Per-Region-OverlapBergmann et al. (2020) to better evaluate model’s capability, denoted by PRO.
Implement Details. Each image is resized to $224\times 224$. The feature extractor is initialized by pre trained DINO(ViT-S-16) Caron et al. (2021), which is frozen during training. We employ AdamWLoshchilov and Hutter (2019) optimizer with a weight decay of $1\times 10^{-4}$. We train our framework for 500 epochs on a single GPU (NVIDIA RTX 4090) with a batch size of 32. The learning rate is initially set to $1\times 10^{-3}$ and reduced by a factor of 0.1 after 200 epochs. The ratio of added noise at the image level linearly increases from 0.6 to 1.0 from the 100th to 400th epochs. The intensity of added noise $\alpha$ at the feature level linearly increases from 0 to 1.0 from the 1st to 400th epochs. The parameter $\eta$ in EMA is set to 0.9999, and the mean teacher is updated every 10 steps. In particular, we performed a 32-fold data augmentation (rotate and flip) under a few-shot setting.
Table 1: Comparison of image-level/pixel-level results on MVTec AD under multi-class setting. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
@@ 4.1 Experimental Setups
Bottle 97.5 / 87.6 / 80.7 97.7 / 91.2 / 90.6 99.7 / 98.1 / 93.2 99.7 / 98.4 / - 100.0 / 98.8 / 94.3 100.0/98.9/96.1 100.0 / 99.1 / 95.0
Cable 57.8 / 71.3 / 40.1 87.6 / 88.1 / 85.4 95.2 / 97.3 / 85.7 94.8 / 96.8 / - 98.5 / 96.2 / 90.2 98.9/96.2/90.3 99.7 / 98.0 / 92.9
Capsule 65.3 / 50.5 / 27.3 78.3 / 89.7 / 84.5 86.9 / 98.5 /89.7 89.0 / 97.1 / - 95.4 / 98.3 / 92.0 94.5/98.5/93.9 93.8 / 98.5 / 91.9
Hazelnut 93.7 / 96.9 / 78.7 99.2 / 95.7 / 87.4 99.8 / 98.1 / 92.9 99.5 / 98.3 / - 99.8 / 99.0 / 95.2 100.0/99.0/95.8 100.0 / 98.9 / 93.6
Metal_nut 72.8 / 62.2 / 66.4 85.1 / 90.9 / 85.2 99.2 / 94.8 / 84.7 99.1 / 97.3 / - 99.7 / 96.4 / 92.4 99.9/96.9/94.0 100.0 / 95.9 / 92.6
Pill 82.2 / 94.4 / 53.9 78.3 / 89.7 / 81.9 93.7 / 95.0 / 94.7 95.7 / 95.7 / - 96.2 / 98.7 / 95.3 97.5/97.6/96.2 97.6 / 98.2 / 96.2
Screw 92.0 / 95.5 / 55.2 45.5 / 93.7 / 84.0 87.5 / 98.3 / 94.9 90.7 / 97.9 / - 91.3 / 99.0 / 93.5 95.9/99.5/97.3 91.8 / 99.0 / 95.5
Toothbrush 90.6 / 97.7 / 68.9 94.7 / 97.5 / 87.4 94.2 / 98.4 / 88.0 99.7 / 99.0 / - 98.9 / 99.1 / 90.9 98.6/99.0/92.4 100.0 / 98.9 / 91.2
Transistor 74.8 / 64.5 / 39.0 82.0 / 86.0 / 83.2 99.8 / 97.9 / 93.3 99.8 / 95.1 / - 98.8 / 93.9 / 76.8 100.0/96.6/88.3 98.7 / 93.8 / 84.3
Zipper 98.8 / 98.3 / 91.9 99.1 / 97.0 / 90.7 95.8 / 96.8 / 91.6 95.1 / 96.2 / - 97.6 / 95.9 / 87.2 99.6/98.4/95.1 99.7 / 97.6 / 90.5
Carpet 98.0 / 98.6 / 93.1 95.9 / 92.4 / 90.6 99.8 / 98.5 / 93.6 99.4 / 98.6 / - 99.5 / 99.0/ 94.7 99.9/99.3/97.5 100.0 / 99.2 / 95.6
Grid 99.3 / 98.7 / 92.1 49.8 / 46.7 / 88.6 98.2 / 96.5 / 91.1 98.5 / 96.6 / - 99.7 / 98.6 /95.8 100.0/99.3/97.3 99.9 / 98.9 / 96.3
Leather 98.7 / 97.3 / 88.5 93.9 / 96.9 / 92.7 100.0 / 98.8 / 96.5 99.8 / 98.8 / - 100.0 / 99.6 / 97.9 100.0/99.4/98.8 100.0 / 97.5 / 97.6
Tile 99.8 / 98.0 / 97.0 93.7 / 93.1 / 90.6 99.3 / 91.8 / 78.6 96.8 / 92.4 / - 100.0 / 96.6 / 87.0 98.6/93.9/81.3 100.0 / 97.4 / 90.2
Wood 99.8 / 96.0 / 94.2 95.2 / 84.8 / 76.3 98.6 / 93.2 / 85.0 99.7 / 93.3 / - 98.7 / 96.4 / 88.0 99.0/94.7/92.4 98.9 / 96.6 / 90.2
Mean 88.1 / 87.2 / 71.1 85.1 / 88.9 / 86.5 96.5 / 96.8 / 90.2 97.2 / 96.8 / 90.7 98.3 / 97.7 / 91.4 98.6/97.7/93.1 98.7 / 98.0 / 92.9
@@
4.2 Anomaly Detection on MVTec-AD
@@
Baseline. We comprehensively evaluated representative methods in both one-class and multi-class settings. In the multi-class setting, we first compare our method with two representative one-class approaches: DRAEM Zavrtanik et al. (2021) and SimpleNet Liu et al. (2023), and then compare with methods specifically designed for the multi-class setting, including UniAD You et al. (2022), RD4AD Deng and Li (2022), DiADHe et al. (2024), ViTAD Zhang et al. (2023), MambaAD He et al. (2024b). Note that all methods are trained under the multi-class scenario. In the one-class setting, the compared models includes: RD4AD Deng and Li (2022), PatchCore Roth et al. (2022), DRAEM Zavrtanik et al. (2021), DeSTSeg Zhang et al. (2023), Simplenet Liu et al. (2023).
@@
Evaluation on MVTec-AD under the multi-class setting. As shown in Table 1, we conducted comprehensive comparative experiments on the MVTec-AD to demonstrate the superiority of our method. It can be observed that applying one-class methods to the multi-class scenario reflects undesirable performance. Furthermore, compared with existing multi-class methods, we outperform all of them, achieving the best 98.7% and 98.0% I-AUC and P-AUC, respectively. ViTAD is a competitive approach, while we outperform it in 12 out of 15 categories in detection capacity, and exceed it in 10 categories for localization performance. In the more challenging PRO metric, our model also achieves a great performance of 92.9%, with a 1.5% improvement over ViTAD. Compared to the MambaAD approach utilizing the latest Mamba-architecture decoder, our method achieves improvements of +0.1% in I-AUC and +0.2% in P-AUC, despite a 0.2% lower performance on the PRO metric.
The qualitative comparison in Figure 3 demonstrates the superior anomaly localization capability of our method compared to DiAD and UniAD. As shown in columns 4 and 10, our approach achieves exceptional precision in handling geometrically complex objects (toothbrushes) and textured surfaces (carpets). The heatmaps which generate by the proposed model accurately align with anomaly contours at the pixel level.
@@
Table 2: Comparison of image-level/pixel-level results on MVTec AD under one-class setting. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
@@
Evaluation on MVTec-AD under the one-class setting. As shown in Table 2, we can observe that compared with the two synthetic data-based methods, DRAEM Zavrtanik et al. (2021) and DeSTSeg Roth et al. (2022), we outperform them by a large margin. Specifically, compared with DRAEM, we outperform it by 1.2% and 1.0% in I-AUC and P-AUC, respectively. And compared to DeSTSeg, we also bring gains of 0.6% I-AUC and 0.4% P-AUC. Moreover, we achieved the best localization result of 98.3% P-AUC, while the detection performance is slightly lower than SimpleNet, ranking second among previous arts. It should be noted that when SimpleNet is adapted to the multi-class setup, its performance significantly drops, as evidenced in Table 1. In contrast, our method demonstrates robust performance across both setups, highlighting its strong generalization capability and suitability for diverse real-world applications.
Table 3: Anomaly detection results on VisA. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
@@
4.3 Anomaly Detection on VisA
Baseline. We further conduct experiments on the challenging VisA dataset. On this dataset, we train a unified model with images of all categories. For comprehensive comparisons, we have selected state-of-the-art models including DRAEM Zavrtanik et al. (2021), UniAD You et al. (2022), Simplenet Liu et al. (2023), ViTAD Zhang et al. (2023), and DiAD He et al. (2024).
Evaluation on VisA. As shown in Table 3, our method once gain achieves outstanding performance. Compared with the pioneering unified framework UniAD, we outperform it greatly by 6.8% and 3.5% in I-AUC and P-AUC, respectively. When compared with the pure ViT based model ViTAD, we outperform it in 10 out of 12 categories for detection and exceed it in 6 categories for localization. Specially, we improve upon it by 1.9%, 0.2% and 0.4% in I-AUC, P-AUC and PRO, respectively. DiAD is one diffusion-based framework for multi-class anomaly detection. We surpass it by 1.5% and 1.2% in I-AUC and P-AUC, respectively. The consistent improvements over the competitors validate the superiority of our method.
4.4 Anomaly Detection on MVTec-3D
@@
Table 4: Anomaly detection results on MVTec-3D. denotes reproduced results using official code. Bold and underline indicate the best and second best, respectively.
Metric
DRÆM*
(ICCV 2021)
UniAD*
(NeurIPS 2022)
DiAD
(AAAI 2024)
ViTAD
(arXiv 2023)
Ours
I-AUC 65.2 78.9 84.6 78.7 84.9
P-AUC 93.2 96.5 96.4 98.0 98.0
PRO 55.0 88.1 87.8 91.3 92.9
Baseline. To further verify the superiority of the proposed method, we conduct experiments on the MVTec-3D dataset and compare it with state-of-the-art models including DRAEM Zavrtanik et al. (2021), UniAD You et al. (2022), and DiAD He et al. (2024).
Evaluation on MVTec-3D. As illustrated in Table 4, our approach outperforms all others across every metric evaluated. Notably, when compared to UniAD, our model demonstrates a considerable advantage, with improvements of 6.0%, 1.5%, and 4.8% in the I-AUC, P-AUC, and PRO, respectively. Compared to the recent diffusion-based method DiAD, our approach has achieved significant improvements in the evaluation metrics, with enhancements of 0.3%, 1.6%, and 5.1% in the I-AUC, P-AUC, and PRO, respectively. ViTAD, a competitive method based on Vision Transformers, also performs well. However, our method still outperforms ViTAD in I-AUC and PRO, further highlighting its effectiveness. These results substantiate the superior performance of our method.
4.5 Ablation Study
Table 5: Ablation studies on attention-guided noise. ‘-’ means no noise is added. ‘R’ indicates adding random noise. ‘A’ denotes adding attention-guided noise.
Noise Type No Noise Only Random Only Attention Hybrid
Image/Feature-level -/- R/- -/R R/R A/- -/A A/A A/R R/A
I-AUC (%) 90.5 90.7 98.0 97.7 91.9 98.3 98.7 98.3 98.6
P-AUC (%) 95.5 95.5 97.7 97.8 96.0 97.7 98.0 97.8 97.8
Attention-Guided Noise. Table 5 studies the impact of noise addition techniques. We can observe that (1) without perturbing the input, this baseline model only achieves inadequate results of 90.5% and 95.5% detection (I-AUC) and localization (P-AUC) capacities, repectively. Such unsatisfactory results suggest that without perturbation the model also well-reconstruct the anomalous patterns. (2) When perturbing the input at the image level in a random manner, we observe a slight increase in I-AUC. However, adding noise at the feature level results in a significant increase, with gains of +7.5% I-AUC and +2.2% P-AUC, respectively. The phenomenon shows that random noise in the image and feature levels forces the model to learn normal patterns more comprehensively, effectively preventing the model from reconstructing anomaly regions. (3) When we introduce noise at the image-/feature-level using our Attention-guided method, compared to randomly adding noise at the same levels, we achieve significant benefits. Specifically, in terms of I-AUC, we bring gains of 1.2% (91.9% vs. 90.7%) and 0.3% (98.3% vs. 98.0%), indicating the effectiveness of our strategy. Moreover, applying our strategy to both image and feature levels further enhances the model’s ability for anomaly detection and localization, achieving the best 98.7% I-AUC and 98.0% P-AUC. (4) Lastly, employing a hybrid noise addition approach only achieves suboptimal performance, further proving the superiority of our attention-guided strategy.
Table 6: Ablation studies on backbones and compare with the previous methods in multi-class setting.
Backbone Method I-AUC/P-AUC(%)
WideResNet-50 SimpleNet 85.1 / 88.9
Ours 87.4 / 93.4
EfficientNet-b4 UniAD 96.5 / 96.8
Ours 97.0 / 96.7
CLIP(ViT-B) ViTAD-CLIP 71.2 / 81.6
Ours 94.6 / 95.8
DINO(ViT-S) ViTAD 98.3 / 97.7
Ours 98.7 / 98.0
Backbone. To prove the generality of our method, we conduct comprehensive experiments using various backbones. As shown in Table 6, We employed two mainstream CNN-based backbones (Wide-Resnet-50He et al. (2015) and EfficientNet-b4Tan and Le (2020)) and two mainstream ViT-based backbones (CLIPRadford et al. (2021) and DINOCaron et al. (2021)) for our experiments. Compared to other mainstream anomaly detection frameworks, our method maintains competitive performance while using the same pre-trained backbones. Specifically, comparing to the ViTAD-CLIP, our method achieves a 23.4% improvement in I-AUC and a 14.2% improvement in P-AUC. Given the reconstruction-based strategy and plain Transformer architecture adopted by ViTAD, this significant enhancement can be attributed to the dynamic difficulty escalation and learnable updates provided by our attention mechanism, which effectively prevents training stagnation and ensures the robustness of anomaly detection. When facing SimpleNet with WideResNet50, our method outperforms with a 2.3% increase in I-AUC and a 4.5% increase in P-AUC. Furthermore, the consistent performance across both CNN and ViT-based architectures suggests that our method is not constrained by the specific characteristics of the backbone models. Instead, it effectively integrates with various feature extraction mechanisms, enhancing their anomaly detection capabilities. This adaptability is crucial for real-world applications where the choice of backbone may vary based on computational constraints or specific task requirements. In conclusion, the results fully demonstrate the generality of the proposed attention-guided perturbation method.
Multi-Layer Features. Table 8 systematically studies the impact of using multi-layer features fusion through four configurations: (1) Using only the deepest layer $f_{11}$achieves suboptimal 97.7% I-AUC and 97.8% P-AUC. This aligns with the fact that high-level features lose spatial details through complex neural networks, limiting precise anomaly location and local anomaly region perception. (2) Compared with only using the deepest feature, combining shallow layers features $f_{2},f_{5}$ boosts detection (98.1% I-AUC) yet degrades localization performance(97.0% P-AUC). We attribute this to low-level noise amplification. While early layers retain object boundaries and detailed information, their sensitivity to texture variations increases false positives in complex backgrounds. (3) Fusing the mid-level layers’ features($f_{8}$) and the deepest layers’ features ($f_{11}$) could improve the performance at image-level, compared with only utilizing the deepest layer features. This further proves that the shallower features have a better representation of the basic texture and contour of the sample to help the model achieve better image-level performance. 4) Overall, integrating both shallow and deep features yields optimal performance. This evidences our core hypothesis: Multi-layer fusion is non-additive but emergent. Just when features from both shallow and deep features interact cooperatively, the model simultaneously resolves different-grained anomaly analyses.
Table 7: Ablation Studies on Noise Intensity $\alpha$ at Feature-level and Attention Map. ‘D’ indicates just using the $A_{\text{learn}}$ to guide the noise tensor, ‘L’ denotes just using the $A_{\text{prior}}$ to guide the noise tensor, and ‘B’ means combining $A_{\text{prior}}$ and $A_{\text{learn}}$ to guide the noise tensor.
Components Noise Intensity $\alpha$ at Feature-level Attention map
Options 0.5 1.0 1.5 0-1.0 L D B
I-AUC (%) 98.5 98.6 98.3 98.7 98.4 98.6 98.7
P-AUC (%) 97.9 97.9 97.9 98.0 97.9 97.9 98.0
Mean Teacher Decoder. Table 8 validates the effectiveness of the Mean Teacher decoder. Using the decoder of the reconstructed model to generate attention maps serves as the baseline. Introducing the Mean Teacher decoder results in an improvement of 0.3% for detection and 0.1% for localization, respectively. The experimental results demonstrate that the Mean Teacher decoder effectively mitigates the high variance of learnable attention weights during the initial training stages. Specifically, the Mean Teacher decoder supports a more stabilized attention-guided perturbation strategy, ensuring more stable weight updates in essential early training phases. This approach contrasts with the baseline decoder, where perturbations in initialization-sensitive regions (such as object boundaries or texture anomalies) induce unstable gradient updates that degrade detection consistency and localization precision.
Easy-to-Hard Noise Addition. As analyzed in Sec. 3.3, the hyperparameter $\alpha$ controls feature-level noise intensity in our noise addition curriculum.Table 7 explores the impact of noise intensity. Specially, we can observe that (1)Low noise intensity ($\alpha=0.5$) reduces the difficulty of decoder reconstruction. It is not conducive to modeling normal samples. This causes the model to fall behind in image-level evaluation (98.5% vs. 98.7%). (2) Excessive noise intensity($\alpha=1.5$)induces severe semantic distortion in feature representations, leading to performance degradation of 0.4% I-AUC and 0.1% P-AUC compared to the optimal configuration. (3) Appropriate noise intensity($\alpha=1$) enhances the model’s ability to capture normal distribution and enables smooth training, resulting in near-optimal performance. (4)The proposed easy-to-hard noise with suitable intensity helps the model capture the target distribution and achieve the best performance. Its advantage is that it dynamically adjusts the noise intensity, which facilitates the model to learn the de-noise task from easy to hard. The results support the advantages of this method.
Table 8: Ablation Studies on Multi-level Features and Mean Distillation. “M.T.D” indicates generating attention maps with mean distillation, while “w.o.” means without using mean distillation.
Components Multi-layer Mean Teacher
Options $f_{11}$ $f_{2},f_{5}$ $f_{8},f_{11}$ $f_{2},f_{5},f_{8},f_{11}$ w.o. M.T.D
I-AUC (%) 97.7 98.1 98.1 98.7 98.4 98.7
P-AUC (%) 97.8 97.0 97.8 98.0 97.9 98.0
Attention Map. Table 7 ablates two components for the final attention mask. We can observe that 1) using $A_{\text{prior}}$ alone to guide the noise tensor already gives us satisfactory performance, and 2) only $A_{\text{learn}}$ guiding the noise tensor can achieve results close to optimal. 3) Finally, combining both obtains the highest performance. In particular, the strong performance achieved with learnable attention mask $A_{\text{learn}}$ further demonstrates the superiority of attention-directed perturbation schemes. Specifically, the $A_{\text{learn}}$ escalates denoising difficulty during normal distribution modeling, implementing efficient dynamic updating that prevents training plateauing in later phases. At the same time, $A_{\text{prior}}$ provides good prior experience in the early stages of training.Their fusion optimally balances exploratory adaptation and exploitation of prior constraints.
4.6 Efficiency Comparison with SoTAs
Table 9: Efficiency comparison of different methods
Method Learnable Parameters(M) FLOPs(G) FPS
DRÆM (ICCV 2021) 97.42 198.15 95.17
UniAD (NeurIPS 2022) 7.48 6.46 286.92
DeSTSeg (CVPR 2023) 32.37 30.67 386.76
SimpleNet (CVPR 2023) 3.94 18.34 63.66
Ours 4.42 5.63 326.65
In addition to model performance, the efficiency of the model is a critical concern for practical applications. We evaluate the efficiency of various methods on a single RTX4090 GPU with a batch size of 32 using Learnable Parameters (M), Floating Point Operations (FLOPs), and Frames Per Second (FPS). As demonstrated in Table 9, our method ranks second in terms of both Learnable Parameters and FPS, while achieving the highest FLOPs. These results indicate that our proposed method effectively balances efficiency and performance, exhibiting strong potential for deployment on end-user devices.
4.7 Extending to Few-shot Anomaly Detection
To further demonstrate the generality and potential of our framework, we conduct experiments on MVTec-AD under the challenging few-shot anomaly detection setups. We compared our method with some competitive methods and the models designed specifically for few-shot setting, including SPADE Cohen and Hoshen (2020), PaDiM Defard et al. (2021), PatchCore Roth et al. (2022), WinCLIP+ Jeong et al. (2023), RWDA Tamura (2023), FastRcon Fang et al. (2023), PromptAD Li et al. (2024). As shown in Table 10, our method also obtains a leading few-shot anomaly detection performance, especially at pixel level. Specifically, our method achieves the best localization performance of 96.2% and 97.3% P-AUC under 2-shot and 4-shot settings. Compared to the competitive method PromptAD, which depends on a large pretrained CLIP model and intricate prompt tuning, we achieved a 0.8% P-AUC increase under 4-shot settings. Patchcore is a competitive anomaly detection model that relies on a resource-intensive memory bank to store nominal patch features. However, when faced with the more challenging few-shot setup, our approach remarkably outperforms it, with improvements of 6.4%, 7.0%, and 9.0% in I-AUC, and 3.0%, 2.9%, and 3.6% in P-AUC under the 4-shot, 2-shot, and 1-shot settings, respectively. The outstanding performance of our framework in few-shot settings highlights the unique advantages of our attention-guided perturbation strategy for scenarios with severely limited data. This success is rooted in the strategy’s ability to prioritize perturbations in semantically critical regions of the input samples, which fundamentally accelerates model convergence and maximizes the utility of every available training sample. By dynamically focusing on areas that are most informative for modeling normal distribution, the attention-guided perturbation method ensures robust normal distribution modeling even with minimal training data, making it suited for few-shot setting.
Table 10: Comparison of image-level/pixel-level results on MVTec under few shot setting. Bold and underline indicate the best and the second best, respectively.